OpenStudy (anonymous):

Hockey teams receive 2 points when they win, and 1 point when they tie. One season a team won a championship with 64 points. They won 11 more games than they tied. How many ties did the team have?

OpenStudy (anonymous):

let the no of games tied = x then no of games won = x + 11 so points scored for x tied games = 1*x = x points scored for (x+11) won games = 2*(x+11) = (2x + 22) ATQ, x + (2x + 22) = 64 ==> 3x + 22 == 64 ==> 3x = 64-22 ==> x = 42/3 = 14 so no of tied games = x = 14 no of won games = x + 11 = 25 so total games played in the season = 14+25 = 39

OpenStudy (anonymous):

if u like the answer, pls click GOOD ANSWER button for me and pls become my fan ( :})